A particle of mass m is projected vertically up with an initial velocity of Vo . If the force due to the friction of the air is directly proportional to its instantaneous velocity,
Calculate the velocity and position of the particle as a function of time?
1
Expert's answer
2020-06-18T11:21:07-0400

As per the question,

Mass of the particle "=M"

Initial velocity of the particle "=V_o"

Let the particle is moving upwards, so force of friction due to air resistance will downwards.

"F=-mg-kv_o"

Where k is constant.

"t=\int_{v_o}^{v}\frac{mdv}{F(v)}"


"=-m\int_{v_o}^{v}\frac{ dv}{mg+kv}"


"=\frac{-m}{k}\ln(mg+kv)|_{vo}^v"


"=\frac{-m}{k}\ln(\frac{mg+kv_o}{mg+kv})"


"\Rightarrow \frac{-kt}{m}=\ln(\frac{mg+kv_o}{mg+kv})"


"\Rightarrow \frac{mg+kv_o}{mg+kv}=e^{\frac{-kt}{m}}"


"\Rightarrow v=(\frac{mg}{k}+v_o)e^{\frac{-k}{m}t}-\frac{mg}{k}"


"v=\frac{dx}{dt}"


Hence,

"dx=((\frac{mg}{k}+v_o)e^{\frac{-k}{m}t}-\frac{mg}{k})dt"


Now, taking the integration of both side of the above equation,

"\int_{x_o}^{x} dx=\int_o^t ((\frac{mg}{k}+v_o)e^{\frac{-k}{m}t}-\frac{mg}{k})dt"


"\Rightarrow x=x_o-\frac{mgt}{k}+(\frac{m^2 g}{k^2}+\frac{mv_o}{k})[1-e^{\frac{-kt}{m}}]"